# FedSA: A Unified Representation Learning via Semantic Anchors for Prototype-based Federated Learning

### v1

@@ Introduction
Fig. LABEL:fig:com shows the accuracy in the cross-silo setting using the HtFE3 and HtFE8 settings under both statistical and model heterogeneity. We observe that the model-based method LG-FedAvg and the KD-based method FedGen show faster convergence compared to prototype-based methods. In the more heterogeneous HtFE8 setting, the convergence curve of prototype-based methods initially exhibits significant volatility, due to the inadequate feature extraction capabilities of heterogeneous models in the early stages of training. Moreover, prototypes are more susceptible to the impacts of statistical and model heterogeneity, resulting in lower accuracy for prototype-based methods. By guiding class representations toward well-separated semantic anchors, FedSA can effectively increase accuracy without slowing down convergence, while also ensuring privacy preservation.
@@ Conclusion
In this paper, we propose a novel framework named FedSA, which introduces simple yet effective semantic anchors to decouple the generation of prototypes from local representation learning. By using these semantic anchors to serve as prototypes, we further propose RMCL and CC to correct local models, enabling them to learn consistent representations and decision boundaries. We then update the semantic anchors with global prototypes via an EMA update to facilitate FL collaboration, which iteratively encourages clients to learn a unified data representation. Extensive experiments and ablation studies under both statistical and model heterogeneity settings demonstrate the superiority of FedSA over prototype-based FL methods.

### v2

@@ Introduction
Figure 3: The test accuracy (%) on Cifar100 in the cross-silo setting using the HtFE8 setting under both statistical and model heterogeneity.
Fig. 3 shows the accuracy in the cross-silo setting using the HtFE8 settings under both statistical and model heterogeneity. We observe that the model-based method LG-FedAvg and the KD-based method FedGen show faster convergence compared to prototype-based methods. The convergence curve of prototype-based methods FedProto and FedTGP initially exhibits significant volatility, due to the inadequate feature extraction capabilities of heterogeneous models in the early stages of training. Moreover, prototypes are more susceptible to the impacts of statistical and model heterogeneity, resulting in lower accuracy for prototype-based methods. By guiding class representations toward well-separated semantic anchors, FedSA can effectively increase accuracy without slowing down convergence, while also ensuring privacy preservation.
@@ Conclusion
In this paper, we propose a novel framework named FedSA, which introduces simple yet effective semantic anchors to decouple the generation of prototypes. By incorporating semantic anchors, we further propose RMCL and CC to correct local models, enabling them to learn consistent representations and decision boundaries. We then update the semantic anchors with global prototypes via an EMA update to encourage clients to learn a unified data representation. Extensive experiments and ablation studies demonstrate the superiority of FedSA over prototype-based FL methods.